OpenStudy (anonymous):

Find the distance between points (-2,-3) and (4,-3)

OpenStudy (anonymous):

Distance between points: sqrt((x2-x1)^2+(y2-y1)^2)

OpenStudy (anonymous):

In this case it's actually x2-x1 = 4+2 = 6